OpenStudy (anonymous):

For element Cl: A) calculate number of miles in 9.36 g.....I did this and got 0.246 Need help with B) calculate the number of grams in 0.104 mil X) how many miles are there in 9.56x10^22 atoms of the element

OpenStudy (ilovebmth1234):

i dont know what question i am supposed to answer

OpenStudy (cuanchi):

B) I told you to use the inverse conversion factor to convert from moles to g and from g to mole 1 mol = 34.5 g 0.140 mol x 34.5 g/ 1 mol = ???? C) for this you need the Avogadro's number 6.022 x 10^23 = 1 mole 9.56x10^22 atoms x 1 mole / (6.02 x 10^23 atoms )= ????? mole

OpenStudy (anonymous):

Okay thank you I wasnt sure how to write the numbers like in what order @Cuanchi